OpenStudy (anonymous):

What is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form? please help me i have 2 minutes to turn this answer in :(

OpenStudy (anonymous):

First find the slope of the perpendicular line. It is the negative reciprocal of the original slope. What is the slope of your starting line?

OpenStudy (anonymous):

thats all it gives me and these are the possible answers: y = –3x – 6 y = negative one-thirdx – 5 y = 3x + 4 y = one-thirdx + 3

OpenStudy (anonymous):

I'm not asking what it gives you. I'm asking you to look at the equation of the line, in the form y = mx+b, where m is the slope and b is the intercept, and tell me what the slope is.

OpenStudy (anonymous):

m is the slope and its -3

OpenStudy (anonymous):

Right, so if m is the slope, the negative reciprocal is \(-\frac{1}{m}\) So in this case we have \(-\frac{1}{-3}\) which simplifies to just?